OpenStudy (anonymous):

Text says the limit of (a^h - 1)/h as h approaches 0 is equal to the derviative of f at 0. I don't understand how this works out. Please explain. Thank you!

ganeshie8 (ganeshie8):

@ikram002p

OpenStudy (ikram002p):

its cuz of LH rule

OpenStudy (ikram002p):

or u wanna use definition to prove it ?

ganeshie8 (ganeshie8):

i think they want to use the limit definition of derivative

ganeshie8 (ganeshie8):

f(x) = a^x

OpenStudy (ikram002p):

i think so as well .

OpenStudy (ikram002p):

its like direct ..

ganeshie8 (ganeshie8):

\[f'(a) = \lim\limits_{h\to a} \dfrac{f(h)-f(a)}{x-a}\]

ganeshie8 (ganeshie8):

substitute f(h) = a^h and a = 0